OpenStudy (anonymous):

Substitution help?? Solve using substitution: 3y − x − 1 = 0 - 4x = -6y Select one: a. No solution b. Infinitely many solutions c. (7, 8/3) d. (1, 2/3)

jimthompson5910 (jim_thompson5910):

I'm going to start off by solving 3y - x - 1 = 0 for x

jimthompson5910 (jim_thompson5910):

3y - x - 1 = 0 3y - x - 1+x = 0+x ... Add x to both sides. 3y - 1 = x x = 3y - 1

jimthompson5910 (jim_thompson5910):

Now that we have isolated x, we can plug that into equation 2 -4x = -6y -4(3y - 1) = -6y ... replace x with 3y - 1 making sense so far?

OpenStudy (anonymous):

yes yes

jimthompson5910 (jim_thompson5910):

are you able to solve -4(3y - 1) = -6y for y?

OpenStudy (anonymous):

9=y right?

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

-4(3y - 1) = -6y -12y + 4 = -6y 4 = -6y + 12y 4 = 6y y = ???

OpenStudy (anonymous):

-2? Sorry I suck at math..

jimthompson5910 (jim_thompson5910):

you divide both sides by 6 to isolate y

OpenStudy (anonymous):

.6

jimthompson5910 (jim_thompson5910):

0.6 repeating, that is equal to 2/3

jimthompson5910 (jim_thompson5910):

so y = 2/3

jimthompson5910 (jim_thompson5910):

then you take this solution for y and plug it into -4x = -6y then solve for x

OpenStudy (anonymous):

awesome bro thanks

jimthompson5910 (jim_thompson5910):

np